24- Problem 7. Evaluate T -?- dy da. 0 0
Answer #1

2 54 x² _x²9² ·dy.da 2- 24 (ega s .. given 4-22 dy) dx The region of integration is. en by y=0; y=N4-82; x=0,x=2 Y=0, x°+9=4712 2 S 12 er drj dio solve for setar dr ier Using? dir del = -28 dr = -1/2 du (dy e -Y? - 4 е heuFrom -2 72 do S Ezerege 7/2 (1) [64-JO مه (-)؟ do 2e4 4. e 7/2 (). 2 e 7 2e (1-0) ale44) 404 is recumed solution .